A 26.0-kg shell is fired from a gun with a muzzle velocity 170.0 m/s at 68.0o above the horizontal. At the top of the trajectory, the shell explodes into two fragments of equal mass. One fragment, whose speed immediately after the explosion is zero, falls vertically. What is the horizontal speed of the other fragment?
Answer #1


momentum before firing


Pix = m*V*costheta

after explosion final momentum Pfx = m/2*vx

from conservation of momentum

Pfx = Pix

m/2*vx = m*v*costheta


26/2*vx = 26*170*cos68

vx = 127.4 m/s